OpenStudy (anonymous):

find the location of the midpoint , M, of CK? C=-19.2 K=-3.4

OpenStudy (anonymous):

Do you mean \(C(-19, 2)\) and \(K(-3, 4)\)?

OpenStudy (anonymous):

no here ill link the problem

OpenStudy (anonymous):

Link doesn't work...

OpenStudy (anonymous):

oh ok ummm one sec

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

\[\frac{-19.2 + (-3.4)}{2} \implies ?\]

OpenStudy (anonymous):

-20.9?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

i did what you did above

OpenStudy (anonymous):

\[\frac{-19.2 + (-3.4)}{2} \implies \frac{-22.6}{2} \implies -11.3\]

OpenStudy (anonymous):

so the answer is -11.3?

OpenStudy (anonymous):

Yes...